OpenStudy (anonymous):

A polynomial has been factored, as shown below: f(x) = (x − 9)(x + 12)(x − 24) What are the zeros of the polynomial? −9, 9, −12, 12, −24, 24 −9, 12, −24 9, −12, 24 9, 12, 24

OpenStudy (anonymous):

anybody know the answer

Vocaloid (vocaloid):

set each part equal to 0, and solve for x (x-9) = 0 (x+12) = 0 (x-24) = 0

OpenStudy (anonymous):

idk how

OpenStudy (decentnabeel):

9,-12,24

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Vocaloid is that correct

OpenStudy (anonymous):

its correct

OpenStudy (anonymous):

dude can u even solve x-9=0?

OpenStudy (anonymous):

i mean all you do is add 9 to both sides

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 4x2 − 12x + 9 16x2 + 24x − 9 4a2 − 10a + 25 36b2 − 24b − 16

OpenStudy (anonymous):

@Vocaloid @DecentNabeel @Icedragon50